the perpendicular distance of P[1,2,3] from the lines  x-6/3 =y-7/2 =z-7/-2  is

Grade:12th pass

1 Answers

Vikas TU
6 years ago
 x-6/3=y-7/2=z-7/-2
x-2 = y-7/2-z-7/2
x-2= y-z-7
x-y+z+5=0
Perpendicular distance from a point (a,b,c)to a line(lx+my+nz+d) is given by formula
|la+mb+nc+d|/√(l^2)+m^2+n^2
=1(1)+(-1)(2)+1(3)+5/√(1^2 )+(-1)^2+(1)^2
=1-2+3+5/√1+1+1
=7/√3